OpenStudy (anonymous):

Simplify 18 - 2[x + (x - 5)].

OpenStudy (nurali):

18-2[2x-5] =18-4x+10 =-4x+28

OpenStudy (anonymous):

it only has this options 8 - 4x 28 - 4x 28 - 2x

OpenStudy (nurali):

28 - 4x or -4x+28 both are same.

OpenStudy (anonymous):

thank you